# ReSAM: Refine, Requery, and Reinforce: Self-Prompting Point-Supervised Segmentation for Remote Sensing Images

### v1

@@ Abstract
Interactive segmentation models such as the Segment Anything Model (SAM) have demonstrated remarkable generalization on natural images, but perform suboptimally on remote sensing imagery (RSI) due to severe domain shift and the scarcity of dense annotations. To address this, we propose a self-prompting, point-supervised framework that adapts SAM to RSIs using only sparse point annotations. Our method employs a Refine–Requery–Reinforce loop, where coarse pseudo-masks are generated from initial points (Refine), improved with self-constructed box prompts (Requery), and embeddings are aligned across iterations to reduce confirmation bias (Reinforce). Without relying on full-mask supervision, our approach progressively enhances SAM’s segmentation quality and domain robustness through self-guided prompt adaptation. We evaluate our proposed method on three RSI benchmark datasets, including WHU, HRSID, and NWPU VHR-10, showing that our method consistently surpasses pretrained SAM and recent point-supervised segmentation methods. Our results demonstrate that self-prompting and semantic alignment provide an efficient path towards scalable, point-level adaptation of foundation segmentation models for remote sensing applications.
@@ 1 Introduction
Annotating high-resolution satellite imagery remains prohibitively expensive; a single 10k×10k image can contain thousands of fine-grained objects. Nevertheless, accurate segmentation is crucial for application such as agriculture management, urban planning and environmental monitoring. Semantic segmentation of high resolution remote sensing images (RSIs) underpins these tasks.
Figure 1: While Vanilla SAM depends on manual prompts (point and box), ReSAM introduces a self-prompting Refine–Requery–Reinforce (R³) loop that progressively refines coarse masks into prompted boxes and generates corresponding pseudo labels, enabling robust point-level adaptation without dense supervision.
Table 1: Pixel Leakage Analysis of SAM Model across Different Datasets. This table reports the percentage of leaked pixels (Overlap regions) on predicted output for each dataset and point-based prompting setup.
@@
. However, training accurate segmentation models typically requires dense pixel-wise annotations, which are extremely costly and time-consuming to collect at a large scale [wei2020hrsid, cheng2016learning, maggiori2017can, DeepGlobe18, ji2018fully, waqas2019isaid]. In contrast, point labels are far cheaper to obtain but inherently incomplete, lacking detailed object boundaries and spatial coverage [chan2025sparse]. The annotation bottleneck is exacerbated by the diverse nature of many satellite sensors, which further complicates the situation. Thus, there is a critical need for methods that can take advantage of point annotation to achieve precise segmentation on various RSI data.
@@
Foundation models have recently reshaped visual understanding by enabling prompt-driven, task-agnostic adaptation across diverse domains. Among them, the recently released Segment Anything Model (SAM) [kirillov2023segment, ravi2024sam] is a foundation model for image segmentation trained on billions of masks across millions of images. SAM’s promptable design (accepting points and boxes) and its impressive zero-shot capabilities on many natural image tasks[roy2023sam, shi2023generalist] make it a promising starting point for segmenting remote sensing images[osco2023segment, yan2023ringmo, luo2024sam]. Early exploration [chen2024rsprompter, ding2024adapting, shan2025ros, yang2025high] have shown that SAM can generalize to aerial and orbital images: for instance, ROS-SAM [shan2025ros] fine-tunes SAM with LoRA, enhances the mask decoder with multi-scale and boundary detail features, and redesigns the training and inference pipeline tailored to remote sensing. RS-Prompter [chen2024rsprompter] learns to generate prompt embeddings tailored for remote sensing categories, allowing SAM to produce semantically meaningful instance masks. However, SAM predicts each mask independently given its prompt, without awareness of other objects. Consequently, it may produce overlapping or fragmented masks in cluttered remote sensing scenes as shown in Fig. 2. Each prediction is locally accurate but globally inconsistent, which poses a fundamental limitation when segmenting densely packed RSI objects. Tab. 1 shows the overlap regions across each instances, highlighting pixels leakage, assuming that each pixels belong to a single object. This information can guide mask refinement.
Hence, despite their methodological advances in adapting SAM to remote sensing domains, these methods remain fully supervised, requiring dense per-pixel segmentation labels. To address this, several methods [zhang2024improving, konwer2025enhancing] have applied a self-training approach utilizing both box and point prompts, but point prompts remain the most efficient option for annotation, especially for densely packed objects in RSIs. Importantly, SAM’s mask decoder is vulnerable to semantic ambiguity when using only point clicks; a single point in a crowded scene can cause multiple nearby objects to merge into one mask.
@@
To address the aforementioned challenges, recent studies [wei2024semantic, wu2024wps, liu2025pointsam] have explored a self-training approach using point prompts as minimal supervision. Among them, PointSAM [liu2025pointsam] advances this idea by introducing negative prompts and instance prototype alignment, demonstrating that point-supervised SAM can substantially outperform the original SAM on remote sensing datasets. However, such methods often rely on maintaining large prototype banks for feature alignment, which is memory-intensive and scales poorly to large datasets. In contrast, we propose ReSAM, a self-prompting, point-supervised framework that converts sparse point input into informative box prompts and iteratively refines predictions through a Refine-Requery-Reinforce loop. This design enhances both instance consistency and domain robustness without requiring dense labels or a large memory load. Box prompts typically yield better results than point prompts, as they provide stronger spatial priors that guide the model towards precise object boundaries, while single points offer only coarse localization cues.
@@
To mitigate such inconsistency and reduce confirmation bias, we introduce a Soft Semantic Alignment (SSA). Instead of relying on static prototype banks, SSA maintains a rolling queue of mask embeddings and enforces a soft cosine-similarity constraint to align instance features across augmented views. This lightweight design encourages semantic consistency while avoiding the memory cost associated with prototype clustering.
@@
  • ReSAM: A point-only self-prompting framework that iteratively converts sparse points into high-quality box prompts via a closed-loop Refine–Requery–Reinforce (R³) strategy, eliminating dense supervision.
  • We introduce a Soft Semantic Alignment (SSA) strategy that aligns mask embeddings via a rolling queue and cosine similarity [xia2015learning], ensuring semantic consistency across iterations while avoiding the heavy memory cost of prototype-based methods.
@@ 2 Related Works
SAM as Backbone Adaptation: Vision foundation models like SAM have shown remarkable generalization due to large-scale pretraining. The original Segment Anything Model [kirillov2023segment] trained on a large dataset and supports flexible prompts such as points and boxes. While SAM performs well on natural images, direct application to remote sensing images (RSIs) is limited by domain-specific characteristics like high resolution, variable scale, and objects’ spatial diversity. To address this, several studies [osco2023segment, ma2024sam, ding2024adapting, liu2024rsps, wang2023samrs, zhang2024mw, zhang2025rsam] have adapted SAM for remote sensing. For example, Zero-to-One-Shot [osco2023segment] explored combining SAM with multimodel prompts, including language guidance, to segment aerial imagery with minimal manual input. SAM-DC [ding2024adapting] uses FastSAM’s [zhao2023fast] encoder to extract powerful features and a convolution on an adapter to focus on change-related ground objects. RS-SAM [zhang2025rsam] incorporates multiscale feature fusion and an encoder adapter tailored for satellite imagery, which improved segmentation accuracy on high-resolution scenes. These works underline that, while SAM is a powerful backbone, careful adaptation is needed for domain shift in remote sensing imagery.
@@
Sparse Point Supervision: Point-level supervision is an efficient strategy to reduce annotation cost in remote sensing, as annotators only click a few points per object [chan2025sparse, bearman2016s, kim2023devil]. However, sparse points do not provide a full object contour, making direct model training challenging. Several methods [wei2024semantic, liu2025pointsam, guo2025discriminatively, tang2022active] have leveraged SAM to generate a pseudo-mask from points. For example, PENet [chan2025sparse], which uses a SAM branch to expand point labels into pseudo-mask via feature-similarity propagation. SAPNet [wei2024semantic] uses Multiple Instance Learning (MIL) to select semantically relevant mask proposals from SAM. PointSAM [liu2025pointsam] developed a self-training SAM that relies solely on point prompts, introducing negative prompt calibration and prototype alignment to correct noisy pseudo-labels. These studies highlight that sparse cues can be effectively leveraged, but careful strategies are needed to compensate for missing contour information.
@@
Figure 3: Overview of ReSAM. Weak and strong views generate pseudo masks and self-prompts to iteratively refine SAM outputs. The pipeline includes Refine (clean instance masks), Requery (self-prompting), and Reinforce (Soft Semantic Alignment) with LoRA adaptation for domain-specific learning.
@@
Self-Training with Embedding Alignment: Self-training leverages pseudo-labels from unlabeled data to improve segmentation performance. In remote sensing, SAM’s zero-shot predictions from point approaches can serve as initial pseudo-labels, but naive self-training may propagate errors. Prior approaches [zhang2021prototypical, das2023weakly, liu2025pointsam] mitigate this with feature alignment and consistency regularization. ProDA [zhang2021prototypical] addresses the problem of noisy pseudo labels and dispersed target features. They utilize consistency across augmentation help from a tighter cluster in feature space, improving adaptation. WDASS [das2023weakly] designed asymmetric alignment to preserve target domain structure rather than forcing all feature into the source domain’s distribution. PointSAM [liu2025pointsam] also utilized feature alignment across the target domain distribution. These methods achieve promising results but lack due to heavy cost of memory intake for feature alignment, limited their access to large target datasets. Our method simplifies this by Soft Semantic Alignment (SSA) strategy that maintains a queue of recent object embedding and enforcing similarity across views. Combined with high-confidence pseudo-masks selection, this ensures that the model learns from reliable signal while limiting error accumulation.
@@ 3.1 Preliminaries
[kirillov2023segment] is a promptable image segmentation model. Given an input image $x\in\mathbb{R}^{H\times W\times 3}$ and a user prompt $p$ (e.g., point or box), SAM predicts a segmentation mask $\hat{M}\in[0,1]^{H\times W}$.
$\hat{M}=\phi_{m}\big(\phi_{i}(x;\theta_{i}),\,\phi_{p}(p;\theta_{p});\theta_{m}\big)$ (1)
where $\phi_{i}$ is SAM’s encoder, a large Vision Transformer (ViT) pretrained on natural images; $\phi_{p}$ is the prompt encoder; and $\phi_{m}$ is the mask decoder. However, SAM’s generalization to remote sensing images (RSIs) is limited due to domain shift.
@@
Each training image $x$ is augmented twice:
$x_{w}=t_{w}(x),\quad x_{s}=t_{s}(x),$ (2)
where $t_{w}$ and $t_{s}$ denote weak and strong transformations. The weak view $x_{w}$ produces pseudo masks, which are then converted into self-prompts used to supervise strong view $x_{s}$:
$\phi_{m}(\phi_{i}(x_{w}),\phi_{p}(p))\approx\phi_{m}(\phi_{i}(x_{s}),\phi_{p}(p^{*})),$ (3)
where $p^{*}$ denotes the automatically generated self-prompt. This weak–strong consistency forms the basis of our self-prompting mechanism.
@@ 3.2 Overview of ReSAM
We propose ReSAM (Refine, Requery, and Reinforce), a self-prompting, point-supervised segmentation framework that adapts SAM to diverse data like RSI’s using only sparse point labels. Unlike existing methods that rely on dense mask supervision, ReSAM introduces a Requery refinement loop and semantic feature alignment to improve segmentation quality and prevent confirmation bias.
@@
  • Refine: Initial masks are generated from point prompts, filtered based on overlap between initial mask instances, and used to derive compact instance regions.
  • Requery: The model automatically generates box prompts from refined regions and re-queries SAM to obtain higher-quality masks.
  • Reinforce: Instance-level features are aligned across training iterations using a Soft Semantic Alignment (SSA) to stabilize pseudo-labels and mitigate drift.
@@ 3.3 Refine: Point-to-Region Initialization
Given a weakly augmented image $I_{w}$, ensuring invariance to simple transformations, and sparse point prompts $P^{+}=\{p_{i}\}$, typically provided by weak supervision, the pretrained SAM generates masks $\hat{M}$:
$\hat{M}=\Phi_{m}(\Phi_{i}(I_{w}),\Phi_{p}(P^{+})).$ (5)
For each instance, given the probabilistic output of an instance segmentation model, each pixel $(i,j)$ is associated with a $K$-dimensional probability vector:
@@
The pixel-wise entropy is computed as:
$H_{ij}=-\sum_{k=1}^{K}\hat{M}_{ij}^{(k)}\log\hat{M}_{ij}^{(k)}.$ (7)
where $H_{ij}$ is the entropy at spatial coordinates $(i,j)$. This quantity captures the model’s uncertainty at each spatial location: low entropy indicates confident predictions, while high entropy indicates ambiguous pixels.
We then threshold the entropy to retain confident pixels:
$\tilde{M}_{ij}^{(k)}=\begin{cases}1,&\text{if }H_{ij}<\epsilon,\\[2.84526pt] 0,&\text{otherwise},\end{cases}\qquad k=1,\dots,K$ (8)
where $\epsilon$ is a tunable threshold controlling the trade-off between precision and recall, and $\tilde{M}_{ij}$ defines confident pixels.
Next, we remove overlapping pixels to keep only pixels belonging to a single instance:
$O_{ij}=\sum_{k=1}^{K}\tilde{M}_{ij}^{(k)}.$ (9)
$M^{\text{ref},(k)}_{ij}=\begin{cases}1,&\text{if }\tilde{M}_{ij}^{(k)}=1\text{ and }O_{ij}=1,\\[2.84526pt] 0,&\text{otherwise},\end{cases}$ (10)
for $k$ = 1,…,$K$.
This guarantees that each pixel contributes to only one instance, preventing cross-object leakage. The resulting masks are clean and instance-specific, suitable as region cues for requerying.

### v2

@@ Abstract
Interactive segmentation models such as the Segment Anything Model (SAM) have demonstrated remarkable generalization on natural images, but it performs suboptimally on remote sensing imagery (RSI) due to severe domain shift and the scarcity of dense annotations. To address this, we propose a self-prompting, point-supervised framework that adapts SAM to RSIs using only sparse point annotations. Our method employs a Refine–Requery–Reinforce loop, where coarse pseudo-masks are generated from initial points (Refine), improved with self-constructed box prompts (Requery), and embeddings are aligned across iterations to reduce confirmation bias (Reinforce). Without relying on full-mask supervision, our approach progressively enhances SAM’s segmentation quality and domain robustness through self-guided prompt adaptation. We evaluate our proposed method on three RSI benchmark datasets, including WHU, HRSID, and NWPU VHR-10, showing that our method consistently surpasses pretrained SAM and recent point-supervised segmentation methods. Our results demonstrate that self-prompting and semantic alignment provide an efficient path towards scalable, point-level adaptation of foundation segmentation models for remote sensing applications. Code is available at https://github.com/MNaseerSubhani/ReSAM.git.
@@ 1 Introduction
Annotating high resolution satellite imagery remains prohibitively expensive; a single 10k×10k image can contain thousands of fine grained objects. Nevertheless, accurate segmentation is crucial for applications such as agriculture management, urban planning, and environmental monitoring. Semantic segmentation of high resolution remote sensing images (RSIs) underpins these tasks.
Figure 1: While Vanilla SAM depends on manual prompts (points and boxes), ReSAM introduces a self-prompting Refine–Requery–Reinforce (R³) loop that progressively refines coarse masks into prompted boxes and generates corresponding pseudo labels, enabling robust point level adaptation without dense supervision.
Table 1: Pixel Leakage Analysis of SAM Model across Different Datasets. This table reports the percentage of leaked pixels (overlap regions) on predicted outputs for each dataset and point-based prompting setup.
@@
However, training accurate segmentation models typically requires dense pixel wise annotations, which are extremely costly and time consuming to collect at a large scale [wei2020hrsid, cheng2016learning, maggiori2017can, DeepGlobe18, ji2018fully, waqas2019isaid]. In contrast, point labels are far cheaper to obtain but inherently incomplete, lacking detailed object boundaries and spatial coverage [chan2025sparse]. The annotation bottleneck is exacerbated by the diverse nature of many satellite sensors, which further complicates the situation. Thus, there is a critical need for methods that can take advantage of point annotation to achieve precise segmentation on various RSI data.
Foundation models have recently reshaped visual understanding by enabling prompt-driven, task-agnostic adaptation across diverse domains. Among them, the recently released Segment Anything Model (SAM) [kirillov2023segment, ravi2024sam] is a foundation model for image segmentation trained on billions of masks across millions of images. SAM’s promptable design (accepting points and boxes) and its impressive zero-shot capabilities on many natural image tasks[roy2023sam, shi2023generalist] make it a promising starting point for segmenting remote sensing images[osco2023segment, yan2023ringmo, luo2024sam].
@@
Early explorations [chen2024rsprompter, ding2024adapting, shan2025ros, yang2025high] have shown that SAM can generalize to aerial and orbital images: for instance, ROS-SAM [shan2025ros] fine-tunes SAM with LoRA, enhances the mask decoder with multi-scale, boundary detail features, and redesigns the training and inference pipeline tailored to remote sensing imagery. RS-Prompter [chen2024rsprompter] learns to generate prompt embeddings for remote sensing categories, allowing SAM to produce semantically meaningful instance masks. However, SAM predicts each mask independently given its prompt, without awareness of other objects. Consequently, it may produce overlapping or fragmented masks in cluttered remote sensing scenes as shown in Fig. 2. Each prediction is locally accurate but globally inconsistent, which poses a fundamental limitation when segmenting densely packed RSI objects. Tab. 1 shows the overlap regions across each instance, highlighting pixel leakage, assuming that each pixel belongs to a single object. This information can guide mask refinement.
@@
Hence, despite their methodological advances in adapting SAM to remote sensing domains, these methods remain fully supervised, requiring dense per-pixel segmentation labels. To address this, several methods [zhang2024improving, konwer2025enhancing] have applied a self-training approach utilizing both box and point prompts, but point prompts remain the most efficient option for annotation, especially for densely packed objects in RSIs. Importantly, SAM’s mask decoder is vulnerable to semantic ambiguity when using only point clicks; a single point in a crowded scene can cause multiple nearby objects to merge into one mask. To address the aforementioned challenges, recent studies [wei2024semantic, wu2024wps, liu2025pointsam] have explored a self-training approach using point prompts as minimal supervision. Among them, PointSAM [liu2025pointsam] advances this idea by introducing instance prototype alignment, demonstrating that point-supervised SAM can substantially outperform the original SAM on remote sensing datasets. However, such methods often rely on maintaining large prototype banks for feature alignment, which is memory intensive and scales poorly to large datasets. In contrast, we propose ReSAM, a self-prompting, point-supervised framework that converts sparse point input into informative box prompts and iteratively refines predictions through a Refine-Requery-Reinforce loop. This design enhances both instance consistency and domain robustness without requiring dense labels or a large memory load. We emphasize that box prompts typically yield better results than point prompts, as they provide stronger spatial priors that guide the model toward precise object boundaries, while single points offer only coarse localization cues.
@@
Furthermore, to mitigate inconsistency and reduce confirmation bias, we introduce a Soft Semantic Alignment (SSA), consists of intuition that objects occurring in similar visual context tend to share semantic structure. For example, boats commonly appear in open sea regions, building appears over ground urban textures, and cars tend to co-occur with road surfaces. This natural regularity implies that boxes regions for different images but depicting similar objects should have similar feature representation. Instead of static feature alignment as prototype banks, SSA maintains a rolling queue of mask embeddings and enforces a soft cosine-similarity constraint to align instance features across augmented views. This lightweight design encourages semantic consistency while avoiding the memory cost associated with prototype clustering, especially for large diverse datasets.
@@
  • ReSAM: A point-only self-prompting framework that iteratively converts sparse points into high quality box prompts via a closed loop Refine–Requery–Reinforce (R³) strategy, eliminating dense supervision.
  • We introduce a Soft Semantic Alignment (SSA) strategy that aligns mask embeddings via a rolling queue and cosine similarity [xia2015learning], ensuring semantic consistency across iterations while avoiding the heavy memory cost of prototype based methods.
@@ 2 Related Works
SAM as Backbone Adaptation: Vision foundation models like SAM have shown remarkable generalization due to large scale pretraining. The original Segment Anything Model [kirillov2023segment] was trained on a large dataset and supports flexible prompts such as points and boxes. While SAM performs well on natural images, direct application to remote sensing images (RSIs) is limited by domain-specific characteristics like high resolution, variable scale, and object spatial diversity. To address this, several studies [osco2023segment, ma2024sam, ding2024adapting, liu2024rsps, wang2023samrs, zhang2024mw, zhang2025rsam] have adapted SAM for remote sensing. For example, Zero-to-One-Shot [osco2023segment] explored combining SAM with multimodel prompts, including language guidance, to segment aerial imagery with minimal manual input. SAM-DC [ding2024adapting] uses FastSAM’s [zhao2023fast] encoder to extract powerful features and a convolution on an adapter to focus on change-related ground objects. RS-SAM [zhang2025rsam] incorporates multiscale feature fusion and an encoder adapter tailored for satellite imagery, which improved segmentation accuracy on high-resolution scenes. These works underline that, while SAM is a powerful backbone, careful adaptation is needed for domain shift in remote sensing imagery.
@@
Sparse Point Supervision: Point-level supervision is an efficient strategy to reduce annotation costs in remote sensing, as annotators only click a few points per object [chan2025sparse, bearman2016s, kim2023devil]. However, sparse points do not provide a full contour of the object, making direct model training challenging. Several methods [wei2024semantic, liu2025pointsam, guo2025discriminatively, tang2022active] have leveraged SAM to generate a pseudo-mask from points. For example, PENet [chan2025sparse], which uses a SAM branch to expand point labels into a pseudo-mask via feature-similarity propagation. SAPNet [wei2024semantic] uses Multiple Instance Learning (MIL) to select semantically relevant mask proposals from SAM. PointSAM [liu2025pointsam] introduces a self-training SAM that relies solely on point prompts, introducing negative prompt calibration and prototype alignment to correct noisy pseudo-labels. These studies highlight that sparse cues can be effectively leveraged, but careful strategies are needed to compensate for missing contour information.
@@
Figure 3: Overview of ReSAM. Weak views $\to$ generate pseudo masks and self-prompts to iteratively refine SAM outputs. The pipeline includes Refine (clean instance masks), Requery (self-prompting), and Reinforce (Soft Semantic Alignment) with LoRA adaptation for domain-specific learning, input with strong view $\to$.
@@
Self-Training with Embedding Alignment: Self-training leverages pseudo-labels from unlabeled data to improve segmentation performance. In remote sensing, SAM’s zero-shot predictions from point approaches can serve as initial pseudo-labels, but naive self-training may propagate errors. Previous approaches [zhang2021prototypical, das2023weakly, liu2025pointsam] mitigate this with feature alignment and consistency regularization. ProDA [zhang2021prototypical] addresses the problem of noisy pseudo labels and dispersed target features. They utilize consistency across augmentation help from a tighter cluster in feature space, improving adaptation. WDASS [das2023weakly] designed asymmetric alignment to preserve target domain structure rather than forcing all features into the source domain distribution. PointSAM [liu2025pointsam] also utilized features alignment across the target domain distribution. These methods achieve promising results but lack due to high memory costs for feature alignment, limited their approach to large target datasets. Our method simplifies this by Soft Semantic Alignment (SSA) strategy that maintains a queue of recent object embedding and enforcing similarity across views. Combined with high-confidence pseudo-masks selection, this ensures that the model learns from reliable signal while limiting error accumulation.
@@ 3.1 Preliminaries
[kirillov2023segment] is a promptable image segmentation model. Given an input image $I\in\mathbb{R}^{H\times W\times 3}$ and a user prompt $p$ (e.g., point or box), SAM predicts a segmentation mask ${M}\in[0,1]^{H\times W}$.
${M}=\phi_{m}\big(\phi_{i}(I;\theta_{i}),\,\phi_{p}(p;\theta_{p});\theta_{m}\big)$ (1)
where $\phi_{i}$ is SAM’s encoder, a large Vision Transformer (ViT) pretrained on natural images; $\phi_{p}$ is the prompt encoder; and $\phi_{m}$ is the mask decoder.
@@
Each training image $I$ is augmented twice:
$I_{w}=t_{w}(I),\quad I_{s}=t_{s}(I),$ (2)
where $t_{w}$ and $t_{s}$ denote weak transformation that applies only small, safe horizontal flip, and strong transformations introduce much more aggressive change that contains color, brightness, contrast, and shadow, while keeping the object shape intact. The weak view $I_{w}$ produces pseudo masks, which are then converted into self-prompts used to supervise strong view $I_{s}$:
$\phi_{m}(\phi_{i}(I_{w}),\phi_{p}(p^{*}))\approx\phi_{m}(\phi_{i}(I_{s}),\phi_{p}(p)),$ (3)
where $p^{*}$ denotes the self generated prompts.
@@ 3.2 Overview of ReSAM
We propose ReSAM (Refine, Requery, and Reinforce), a self-prompting, point-supervised segmentation framework that adapts SAM to diverse data such as RSI using only sparse point labels.
@@
  • Refine: Initial masks are generated from point prompts, filtered based on overlap between mask instances, and used to derive compact instance regions.
  • Requery: The model automatically generates box prompts from these refined regions and re-queries SAM to obtain higher-quality masks.
  • Reinforce: Instance level features are aligned across training iterations using Soft Semantic Alignment (SSA) to stabilize pseudo labels and mitigate drift.
@@ 3.3 Refine: Point-to-Region Initialization
Given a weakly augmented image $I_{w}$, ensuring invariance to simple transformations, and sparse point prompts $P^{+}=\{p_{i}\}$, typically provided by weak supervision, the pretrained SAM generates masks $\hat{M}^{(k)}$:
$\hat{M}^{(k)}=\Phi_{m}(\Phi_{i}(I_{w}),\Phi_{p}(P^{+})).$ (5)
For each instance $k$, given the probabilistic output of an instance segmentation model, each pixel $(i,j)$ is associated with a $K$ dimensional probability vector:
@@
We then calculate the Shannon entropy map for each instance:
$H_{ij}^{(k)}=-[\hat{M}_{ij}^{(k)}\log(\hat{M}_{ij}^{(k)})+\big(1-\hat{M}_{ij}^{(k)})\log(1-\hat{M}_{ij}^{(k)})]$ (7)
where $H_{ij}^{(k)}$ is the entropy at spatial coordinates $(i,j)$ for instance $k$, which is normalized to $[0,1]$. This quantity captures the model’s uncertainty at each spatial location: low entropy indicates confident predictions, while high entropy indicates ambiguous pixels.
We then filter out most confident pixel using:
${C}_{ij}^{(k)}=\begin{cases}1,&\text{if }\hat{M}_{ij}^{(k)}(1-H_{ij}^{(k)})>\epsilon,\\[2.84526pt] 0,&\text{otherwise},\end{cases}\qquad k=1,\dots,K$ (8)
where $\epsilon$ is a threshold controlling the trade off between precision and recall, and $\hat{M}_{ij}$ defines confident pixels. We set $\epsilon$ to 0.5.
Next, we remove overlapping pixels so that each pixel belongs to a single instance:
$O_{ij}=\begin{cases}1,&\text{if }\sum_{k=1}^{K}C_{ij}^{(k)}>1,\\[2.84526pt] 0,&\text{otherwise},\end{cases}$ (9)
$M^{\text{ref},(k)}_{ij}={C}_{ij}^{(k)}(1-O_{ij})$ (10)
for $k$ = 1…$K$.
Where $M^{\text{ref},(k)}_{ij}$ refined mask for instance $k$ and $O$ is overlap pixel map. This guarantees that each pixel contributes to only one instance, preventing cross-object leakage. The resulting masks are clean and instance-specific, suitable as region cues for requerying.